OpenStudy (anonymous):

Just wondering about the concepts shouldn't you be able to use greens theorem to derive the polar conversion for double integrals?

OpenStudy (anonymous):

Just wondering if you could derive the \[\int\limits_{\alpha}^{\beta}\int\limits_{a}^{b}f(rcos \theta, rsin \theta)rdrd \theta\]

OpenStudy (anonymous):

without using the summation notation

OpenStudy (anonymous):

There should be some way to do it with greens / a form of line integrals no?